OpenStudy (kaylala):

Trigonometric Equation tanA cosA + √(3) cosA = 0

OpenStudy (anonymous):

sin\[\frac{ sinA }{ CosA } * cosA + \sqrt{3} cosA = 0\] \[sinA+\sqrt{3}cosA = 0\]

OpenStudy (anonymous):

\[\frac{ 1 }{2 }\sin A+\frac{ \sqrt{3} }{2 }\cos A=0\] \[\sin 30\sin A+\cos 30 \cos A=0 or \cos (A-30)=0=\cos 90,\cos 270\] A-30=90,270 A=120,300 degrees.

OpenStudy (anonymous):

I was wrong, missed a factor after dividing.